a. A work center at the Mechanical Lyold Company has 4 machines and each machine is operated 8.5 hours a day, 6 days a week. Last week, the work center operated for a total 158 hours and produced 140 hours of work.

Calculate the
i. The efficiency of the work center
ii. The utilization of the work center

iii. what is the rated capacity of the workcenter
Answer #1

(i) Efficiency = Time of operation/ Time available = 158 8.5 6* 4

Efficiency = 0.7745 = 77.45%

(ii) Utilization = Total Output/ Time operated = 140/158= 0.8861 = 88.61%

(iii) Rated capacity = 8.5*6*4 = 204 hours